The frequency of the transverse waves with an amplitude of 0.02 m moving along a + X axis with a speed of 10 m / s on a string is 5 Hz. Find the wavelength (λ) of the transverse waves traveling on this string and write the displacement equation of these waves. (Take = 3)
1
Expert's answer
2020-06-09T13:21:30-0400

Wave is moving along a positive direction of X-axis.

Amplitude of the wave is "A=2\times 10^{-2}m" .

Speed of wave is "v=10m\/s" ,Frequency is "\nu=5Hz"

Let, wave length of the given transverse wave is "\lambda"


We know that

"v=\nu\lambda"

Thus, on plugin the value we get

"\lambda=\frac{10}{5}=2m"

Wave equation is given by

"y=A\sin(kx-\omega t)\\\implies y=A\sin\big(\frac{2\pi}{\lambda}x-2\pi\nu t\big)"

Hence, on plugin the value we get our wave equation is

"y=2\times 10^{-2}\sin(\pi x-10\pi t)"